# Another look at inference after prediction

### v1

@@ 1 Introduction
Here we build on the work on Motwani and Witten to study the statistical efficiency of the PPI method. We highlight that a more efficient estimator can be obtained by simply adding a weight to the PPI estimator, an approach that was proposed by Chen & Chen (CC) more than 25 years ago in the context of double sampling designs (Chen and Chen, 2000). We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2. We then study the asymptotic properties of the PPI estimator and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. We also point to earlier literature from the statistics and economics literature dating back to the 1960s that provides important historical context for the PB inference problem in Section 4. In Section 5, we evaluate our findings with extensive simulation studies. Our results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice. We close with additional discussion in Section 6.
@@ 2.4 PB inference methods
Motwani and Witten also study the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions. Stated otherwise, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency or to yield reliable hypothesis testing and interval estimation. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al. (1994), and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen (2000); Tsiatis (2006); Tian et al. (2012); Wang and Wang (2015)). We return to this point in Section 4.
@@ 3.1 Asymptotic analysis of the PPI estimator
More specifically, the sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al. (2023a) and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
@@
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
The improvement of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is therefore stronger with increasing correlation of residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$. Therefore, $\widehat{Y}$ must be a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though it does not necessarily need to be derived from a correctly specified ML model. In particular, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency if and only if the correlation of residuals is at least half of the ratio of residual standard deviations of $\widehat{Y}$ to that of $Y$.
@@ 3.2 Asymptotic analysis of the CC estimator
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen (2000) proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology and notation (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is identical to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an augmented estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a simple weight function in front of the augmentation term. In fact, the framework of Robins et al. (1994) implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results provided for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ are detailed in Appendix C.
@@
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of the $\widehat{Y}$ to scale the covariance according to the precision of $\widehat{Y}$. This scaling essentially enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be as least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not possess this property.
@@ 4 Related literature
Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error (Angelopoulos et al., 2023a, b; Gan and Liang, 2023; Miao et al., 2023; Zrnic and Candès, 2024a, b). Here we draw connections between PB inference methods and methods from the economics and statistics literature dating back to the 1960s. These connections reveal that aspects of the PB inference problem have been explored in various areas, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, future research in PB inference could benefit from a more comprehensive review of these and other related approaches than what is provided here.
@@ 4.1 Seemingly unrelated regression & its use in statistical genetics
The SUR estimator was recently applied for a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment (McCaw et al., 2024). The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on the SynSurr estimator, Miao et al. (2024) proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
@@ 4.2 Augmented estimation for quantifying a treatment effect
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ for $k=0,1$ are arbitrary scalar functions of ${\bf{Z}}$. As $X\perp{\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. In the context of PB inference, pre-trained models could be used for predicting $Y$ on the basis of ${\bf{Z}}$ for $h^{(k)}({\bf{Z}})$ for $k=0,1$.
@@ 4.3 Improving precision of Monte Carlo simulations with control variates
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ Rothery (1982); Davidson and MacKinnon (1992). The augmentation term reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later used for mean estimation in the semi-supervised setting by Zhang et al. (2019).
@@ 5.1 Data generation
We perform several simulation studies to compare the performance of the PPI and CC methods. In all settings, we generate two datasets: (i) a training dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ and an unlabeled dataset with realizations of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$. We use 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed. The size of the training set is fixed at 10,000 and the size of the analysis dataset is also set to $n=10,000$. We let $n_{\text{\tiny{lab}}}=1000,2000,3000,4000,5000$ and $n_{\text{\tiny{unlab}}}=n-n_{\text{\tiny{lab}}}$. For both the training and analysis datasets, $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ are generated from a multivariate normal distribution. Each element of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ has a mean of 0, variance of 1, and covariance of 0.4 with the remaining variables. The goal is to conduct inference on the marginal association of $Y$ and $X_{1}$ in a linear model. The parameter of interest is therefore $\beta^{*}_{1}$ defined as
@@
We consider 2 primary scenarios for evaluating the PB inference methods. In both scenarios, the outcome is generated as
$Y=X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{{\sf\scriptscriptstyle{T}}}\bm{\theta}+\epsilon$
where ${\bf{W}}^{{\sf\scriptscriptstyle{T}}}=(W_{1},\dots,W_{q})$ and the error term $\epsilon$ is simulated independently from a standard normal distribution. We varied the coefficients of $\bm{\theta}$ and $\theta_{\text{\tiny{INT}}}$ as well as $q$ and the composition of ${\bf{Z}}$.
@@
We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds. In both sub-scenarios, we let $p=10$, $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0$. In Scenario 1a, we obtain a strong prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{1},\dots,W_{10})^{{\sf\scriptscriptstyle{T}}}$ while in Scenario 1b we obtain a weaker prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{5},X_{1}W_{5})^{{\sf\scriptscriptstyle{T}}}$.
@@
We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$ (e.g., in a genome-wide association study). In Scenario 2a, we let $p=9$, $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},\dots,W_{9})^{{\sf\scriptscriptstyle{T}}}$. In Scenario 2b, we let $p=3$, $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0.25$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},W_{2},W_{3},W_{2}W_{1},W_{3}W_{1},W_{3}W_{2})^{{\sf\scriptscriptstyle{T}}}$. Scenario 2b uses a weaker prediction model than Scenario 2a.
We evaluate the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$. We also include two comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, serving as the baseline, and $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, highlighting the limitations of naive inference. We present the percent bias for each estimator. For the PPI and CC methods, we present the confidence interval (CI) length ratio, computed as the mean of the ratio of the length of the PB inference estimator’s normal-based CI relative to the length of the normal-based CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ (which is simply the ratio of the asymptotic standard errors). The CI length ratio has a baseline value of 1, which indicates that the CI of the PB inference method is the same length as the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. A CI length ratio less than 1 indicates that the PB inference estimator has lower standard error than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.

### v2

@@ 1 Introduction
Here we build on the work on Motwani and Witten to study the statistical efficiency of the PPI method. We show that a more efficient estimator can be obtained by simply adding a weight to the PPI estimator, an approach that was proposed by Chen & Chen (CC) more than 25 years ago in the context of double sampling designs (Chen and Chen, 2000). We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2. We then study the asymptotic properties of the PPI estimator and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. We also point to earlier literature from the statistics and economics literature dating back to the 1960s that provides important historical context for the PB inference problem in Section 4. In Section 5, we evaluate our findings with extensive simulation studies. Our results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice. We close with additional discussion in Section 6.
@@ 2.4 PB inference methods
Motwani and Witten also studied the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions. Stated otherwise, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency or to yield reliable hypothesis testing and interval estimation. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al. (1994), and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen (2000); Tsiatis (2006); Tian et al. (2012); Wang and Wang (2015)). We return to this point in Section 4.
@@ 3.1 Asymptotic analysis of the PPI estimator
The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al. (2023a) and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
@@
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ therefore gains efficiency if and only if the correlation of the residuals is at least half of the ratio of residual standard deviations of $\widehat{Y}$ to that of $Y$. Put another way, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with respect to efficiency when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function. The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
@@ 3.2 Asymptotic analysis of the CC estimator
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen (2000) proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology and notation (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an augmented estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term. In fact, the framework of Robins et al. (1994) implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results provided for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ are detailed in Appendix C.
@@
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\boldsymbol{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
@@ 4 Related literature
Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error (Angelopoulos et al., 2023a, b; Miao et al., 2023; Zrnic and Candès, 2024a, b). Here we draw connections between PB inference methods and methods from the economics and statistics literature dating back to the 1960s. These connections reveal that aspects of the PB inference problem have been explored in various areas, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, future research in PB inference could benefit from a more comprehensive review of these and other related approaches than what is provided here.
@@ 4.1 Seemingly unrelated regression & its use in statistical genetics
The SUR estimator was recently applied for a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment (McCaw et al., 2024). The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on SynSurr, Miao et al. (2024) proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
@@ 4.2 Augmented estimation for quantifying a treatment effect
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of ${\bf{Z}}$ for $k=0,1$. As $X$ and ${\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. In the context of PB inference, pre-trained models could be used for predicting $Y$ on the basis of ${\bf{Z}}$ in the augmentation term.
@@ 4.3 Improving precision of Monte Carlo simulations with control variates
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ (Rothery, 1982; Davidson and MacKinnon, 1992). The augmentation term reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later used for mean estimation in the semi-supervised setting by Zhang et al. (2019).
@@ 5.1 Data generation
We perform several simulation studies to compare the performance of the PPI and CC methods. In all settings, we generate two datasets: (i) a training dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ and an unlabeled dataset with realizations of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$. We use 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed. The size of the training set is set to 10,000 and the size of the analysis dataset is also $n=10,000$. We let $n_{\text{\tiny{lab}}}=1000,2000,3000,4000,5000$ and $n_{\text{\tiny{unlab}}}=n-n_{\text{\tiny{lab}}}$. For both the training and analysis datasets, $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ are generated from a multivariate normal distribution. Each element of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ has a mean of 0, variance of 1, and covariance of 0.4 with the remaining variables. The goal is to conduct inference on the marginal association of $Y$ and $X_{1}$ in a linear model. The parameter of interest is therefore $\beta^{*}_{1}$ defined as
@@
We consider two primary scenarios for evaluating the PB inference methods. In both scenarios, the outcome is generated as
$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{{\sf\scriptscriptstyle{T}}}\bm{\theta}+\epsilon$
where ${\bf{W}}^{{\sf\scriptscriptstyle{T}}}=(W_{1},\dots,W_{q})$ and the error term $\epsilon$ is simulated independently from a standard normal distribution. We varied the coefficients of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$ as well as $q$ and the composition of ${\bf{Z}}$.
@@
We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds. In both sub-scenarios, we let $q=10$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$. In Scenario 1a, we obtain a strong prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{1},\dots,W_{10})^{{\sf\scriptscriptstyle{T}}}$ while in Scenario 1b we obtain a weaker prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{5},X_{1}W_{5})^{{\sf\scriptscriptstyle{T}}}$. The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.9 and 0.6 in Scenarios 1a and 1b, respectively.
@@
We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$. In Scenario 2a, we let $q=9$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},\dots,W_{9})^{{\sf\scriptscriptstyle{T}}}$. In Scenario 2b, we let $q=3$, $\theta_{X_{1}}=2$, $\theta_{\text{\tiny{INT}}}=0.25$, and $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},W_{2},W_{3},W_{2}W_{1},W_{3}W_{1},W_{3}W_{2})^{{\sf\scriptscriptstyle{T}}}$. The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.5 and 0.4 in Scenarios 2a and 2b, respectively.
We evaluate the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$. We also include two comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, serving as the baseline, and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, highlighting the limitations of naive inference. We present the percent bias for each estimator. For the PPI and CC methods, we present the confidence interval (CI) length ratio, computed as the mean of the ratio of the length of the PB inference estimator’s normal-based CI relative to the length of the normal-based CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ (which is simply the ratio of the asymptotic standard errors). The CI length ratio has a baseline value of 1, which indicates that the CI of the PB inference method is the same length as the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. A CI length ratio less than 1 indicates that the PB inference estimator has lower standard error than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.